# Asymptotics for conformal inference

### v1

@@ C.2 Prediction setting
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\left({F^{w}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\left(G^{w}\right)\\ {\rho^{w}}(G^{w})^{\prime}\left[\mathbb{V}\left(I^{w}\right)+\left[I-I^{w}\right]N\right]\end{pmatrix}$
[…]
@@
$\sqrt{m}\left({\widehat{F}}_{m,\tiny\mbox{test}}-{F_{\tiny\mbox{test,cal}}}\right)\overset{\mathcal{L}}{\rightarrow}\mathbb{U}(1-{F_{\tiny\mbox{test,cal}}})\text{ on $D(0,1)$},$
@@
$\sqrt{n}\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}{\rho^{w}}\left(\mathbb{U}\left({{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right)+\left[I-{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}\right]N\right)$
[…]
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}-{F_{\tiny\mbox{test,cal}}}\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\left({F^{w}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\left(1-{F_{\tiny\mbox{test,cal}}}\right)\\ {\rho^{w}}(G^{w})^{\prime}\left[\mathbb{V}\left(I^{w}\right)+\left[I-I^{w}\right]N\right]\end{pmatrix}$
[…]
@@
$\begin{pmatrix}\sqrt{m}\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)\right]\\ \sqrt{n}\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\left({F^{w}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-I)\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\left(1-{F_{\tiny\mbox{test,cal}}}\circ{{F^{w}_{{\tiny\mbox{cal}}}}}{-1}(1-I)\right)\\ {\rho^{w}}(G^{w})^{\prime}\left[\mathbb{V}\left(I^{w}\right)+\left[I-I^{w}\right]N\right]\end{pmatrix}$
[…]
@@ C.3 Novelty detection setting
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{w^{*},(n)}-I\\ {\widehat{{G}}}_{m,{1}}^{w^{*},(n)}-G^{w^{*}}\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+{\rho^{w^{*}}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w^{*}}}+\left[I-I^{w^{*}}\right]N\right)\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G^{w^{*}}}+{(G^{w^{*}})}^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w^{*}}}+\left[I-I^{w^{*}}\right]N\right)\end{pmatrix}$
$\eqqcolon\begin{pmatrix}\mathbb{Z}^{w^{*}}_{0}\\ \mathbb{Z}^{w^{*}}_{1}\end{pmatrix}\text{ on $\left[D(0,1)\right]^{2}$}.$
@@
$\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{w^{*},(n)}-{G_{{\tiny\mbox{mixt}}}^{w^{*}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\left(1-\pi_{0}\right)\sigma^{2}}\mathbb{V}_{G^{w^{*}}}$
$+({{G_{{\tiny\mbox{mixt}}}^{w^{*}}}})^{\prime}{\rho^{w^{*}}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w^{*}}}+\left[I-I^{w^{*}}\right]N\right)$
$=\pi_{0}\mathbb{Z}^{w^{*}}_{0}+\left(1-\pi_{0}\right)\mathbb{Z}^{w^{*}}_{1}\eqqcolon\mathbb{Z}^{w^{*}}\text{ on $D(0,1)$}.$
@@
$\begin{pmatrix}\sqrt{m_{0}(m)}\left({\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)-{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)\right)\\ \sqrt{m_{1}(m)}\left({\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)\right)\\ \sqrt{n}\left({F^{w^{*}}_{{\tiny\mbox{cal}}}}\circ{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)-(1-I)\right)\\ \sqrt{n}\left(F_{\tiny\mbox{test}}\circ{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}\right)^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\left({F^{w^{*}}_{{\tiny\mbox{cal}}}}\right)^{-1}}(1-I)\right)\end{pmatrix}$
@@
$\begin{pmatrix}\mathbb{U}\\ \mathbb{V}\circ G^{w^{*}}\\ {\rho^{w^{*}}}\left(\mathbb{W}\left(I^{w^{*}}\right)+\left[I-I^{w^{*}}\right]N\right)\\ {\rho^{w^{*}}}(G^{w^{*}})^{\prime}\left(\mathbb{W}\left(I^{w^{*}}\right)+\left[I-I^{w^{*}}\right]N\right)\end{pmatrix}\text{ on $\left[D(0,1)\right]^{4}$}.$
@@
$\sqrt{\tau_{n,m}}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})-\mathcal{T}^{{w^{*}}}_{\alpha}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}({\mathcal{T}^{{w^{*}}}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})\right)-\pi_{0}\alpha\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{{G_{{\tiny\mbox{mixt}}}^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}_{0}({\mathcal{T}^{{w^{*}}}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{TDP}_{m}^{(n)}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})\right)-G^{w}(\mathcal{T}^{{w^{*}}}_{\alpha})\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{{(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}({\mathcal{T}^{w}_{\alpha}})+\mathbb{Z}^{w^{*}}_{1}(\mathcal{T}^{w}_{\alpha}),$
@@ Appendix D Weighted novelty detection for a general weight function
$\zeta_{\alpha}(t)$ $=1-(1-\pi_{0})\frac{({G_{0}^{w}})^{\prime}G^{w}-{G_{0}^{w}}({G^{w}})^{\prime}}{\left(\alpha^{-1}-({G_{{\tiny\mbox{mixt}}}^{w}})^{\prime}(t)\right)G_{0}^{w}}$
where $\mathbf{p}(t)$ corresponds to the positive false discovery rate at $t$ (Storey,, 2002). Note that in the oracle case $\zeta_{\alpha}(\mathcal{T}^{w^{*}}_{\alpha})=0$, but $\zeta_{\alpha}(\mathcal{T}^{w}_{\alpha})$ is not necessarily zero for a general $w$.
@@
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{w,(n)}-G_{0}^{w}\\ {\widehat{{G}}}_{m,{1}}^{w,(n)}-G^{w}\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}_{G_{0}^{w}}+({G_{0}^{w}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w}}+\left[I-I^{w}\right]N\right)\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G^{w}}+({G^{w}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w}}+\left[I-I^{w}\right]N\right)\end{pmatrix}\eqqcolon\begin{pmatrix}\mathbb{Z}^{w}_{0}\\ \mathbb{Z}^{w}_{1}\end{pmatrix}$
[…]
@@
$\sqrt{\tau_{n,m}}\left({\widehat{G}}_{m}^{w,(n)}-{G_{{\tiny\mbox{mixt}}}^{w}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}_{G_{0}^{w}}+\sqrt{\left(1-\pi_{0}\right)\sigma^{2}}\mathbb{V}_{G^{w}}$
$+({{G_{{\tiny\mbox{mixt}}}^{w}}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\left(\mathbb{W}_{I^{w}}+\left[I-I^{w}\right]N\right)$
$=\pi_{0}\mathbb{Z}^{w}_{0}+\left(1-\pi_{0}\right)\mathbb{Z}^{w}_{1}\eqqcolon\mathbb{Z}^{w}\text{ on $D(0,1)$}.$
@@
$\sqrt{\tau_{n,m}}\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w,(n)})-\mathcal{T}^{w}_{\alpha}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{w}_{\alpha})}\mathbb{Z}^{w}({\mathcal{T}^{w}_{\alpha}});$
$\sqrt{\tau_{n,m}}\left(\mathrm{FDP}_{m}^{(n)}\left({\mathcal{R}}^{w}_{\alpha}\right)-\mathbf{p}\left(\mathcal{T}^{w}_{\alpha}\right)\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\mathbb{Z}^{w}_{0}({\mathcal{T}^{w}_{\alpha}})}{G_{0}^{w}\left(\mathcal{T}^{w}_{\alpha}\right)}\mathbf{p}\left(\mathcal{T}^{w}_{\alpha}\right)\left[1-\mathbf{p}\left(\mathcal{T}^{w}_{\alpha}\right)\zeta_{\alpha}\left(\mathcal{T}^{w}_{\alpha}\right)\right]$

### v3

@@ C.2 Prediction setting
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({F^{w}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\mathopen{}\mathclose{{\left(G^{w}}}\right)\\ {\rho^{w}}(G^{w})^{\prime}\mathopen{}\mathclose{{\left[\mathbb{V}\circ{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right]\end{pmatrix}$
$\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$},$
@@
$\sqrt{m}\mathopen{}\mathclose{{\left({\widehat{F}}_{m,\tiny\mbox{test}}-{F_{\tiny\mbox{test,cal}}}}}\right)\overset{\mathcal{L}}{\rightarrow}\mathbb{U}(1-{F_{\tiny\mbox{test,cal}}})\text{ on $D(0,1)$},$
@@
$\sqrt{n}\mathopen{}\mathclose{{\left({({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}-({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right)\overset{\mathcal{L}}{\rightarrow}(({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1})^{\prime}{\rho^{w}}\mathopen{}\mathclose{{\left(\mathbb{U}\mathopen{}\mathclose{{\left({{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right)+\mathopen{}\mathclose{{\left[I-{{V}^{w}_{{\tiny\mbox{cal}}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}}}\right]N}}\right)$
$\text{ on $D(0,1)$},$
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}-{F_{\tiny\mbox{test,cal}}}}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({F^{w}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\mathopen{}\mathclose{{\left(1-{F_{\tiny\mbox{test,cal}}}}}\right)\\ {\rho^{w}}(G^{w})^{\prime}\mathopen{}\mathclose{{\left[\mathbb{V}\circ{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right]\end{pmatrix}$
$\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$},$
@@
$\begin{pmatrix}\sqrt{m}\mathopen{}\mathclose{{\left[{\widehat{F}}_{m,\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)}}\right]\\ \sqrt{n}\mathopen{}\mathclose{{\left[F_{\tiny\mbox{test}}\circ{({{\widehat{F}}_{\tiny\mbox{cal}}^{w,(n)}})^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({F^{w}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-I)}}\right]\end{pmatrix}\overset{\mathcal{L}}{\rightarrow}\begin{pmatrix}\mathbb{U}\circ\mathopen{}\mathclose{{\left(1-{F_{\tiny\mbox{test,cal}}}\circ({{F^{w}_{{\tiny\mbox{cal}}}}})^{-1}(1-I)}}\right)\\ {\rho^{w}}(G^{w})^{\prime}\mathopen{}\mathclose{{\left[\mathbb{V}\circ{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right]\end{pmatrix}$
$\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$},$
@@ C.3 Novelty detection setting
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{w^{*},(n)}-I\\ {\widehat{{G}}}_{m,{1}}^{w^{*},(n)}-G^{w^{*}}\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}+{\rho^{w^{*}}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w^{*}}}+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G^{w^{*}}}+{(G^{w^{*}})}^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w^{*}}}+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)\end{pmatrix}$
$\eqqcolon\begin{pmatrix}\mathbb{Z}^{w^{*}}_{0}\\ \mathbb{Z}^{w^{*}}_{1}\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$}.$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{w^{*},(n)}-{G_{{\tiny\mbox{mixt}}}^{w^{*}}}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}+\sqrt{\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\sigma^{2}}\mathbb{V}_{G^{w^{*}}}$
$+({{G_{{\tiny\mbox{mixt}}}^{w^{*}}}})^{\prime}{\rho^{w^{*}}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w^{*}}}+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)$
$=\pi_{0}\mathbb{Z}^{w^{*}}_{0}+\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\mathbb{Z}^{w^{*}}_{1}\eqqcolon\mathbb{Z}^{w^{*}}\text{ on $D(0,1)$}.$
@@
$\begin{pmatrix}\sqrt{m_{0}(m)}\mathopen{}\mathclose{{\left({\widehat{F}}_{{\tiny\mbox{$m$,$0$,test}}}\circ{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)-{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)}}\right)\\ \sqrt{m_{1}(m)}\mathopen{}\mathclose{{\left({\widehat{F}}_{{\tiny\mbox{$m$,$1$,test}}}\circ{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)}}\right)\\ \sqrt{n}\mathopen{}\mathclose{{\left({F^{w^{*}}_{{\tiny\mbox{cal}}}}\circ{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)-(1-I)}}\right)\\ \sqrt{n}\mathopen{}\mathclose{{\left(F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({\widehat{F}}_{\tiny\mbox{cal}}^{{w^{*}},(n)}}}\right)^{-1}}(1-I)-F_{\tiny\mbox{test}}\circ{\mathopen{}\mathclose{{\left({F^{w^{*}}_{{\tiny\mbox{cal}}}}}}\right)^{-1}}(1-I)}}\right)\end{pmatrix}$
@@
$\begin{pmatrix}\mathbb{U}\\ \mathbb{V}\circ G^{w^{*}}\\ {\rho^{w^{*}}}\mathopen{}\mathclose{{\left(\mathbb{W}\mathopen{}\mathclose{{\left(I^{w^{*}}}}\right)+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)\\ {\rho^{w^{*}}}(G^{w^{*}})^{\prime}\mathopen{}\mathclose{{\left(\mathbb{W}\mathopen{}\mathclose{{\left(I^{w^{*}}}}\right)+\mathopen{}\mathclose{{\left[I-I^{w^{*}}}}\right]N}}\right)\end{pmatrix}\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{4}$}.$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})-\mathcal{T}^{{w^{*}}}_{\alpha}}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}({\mathcal{T}^{{w^{*}}}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})}}\right)-\pi_{0}\alpha}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\pi_{0}}{{G_{{\tiny\mbox{mixt}}}^{w^{*}}}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}_{0}({\mathcal{T}^{{w^{*}}}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{TDP}_{m}^{(n)}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w^{*},(n)})}}\right)-G^{w}(\mathcal{T}^{{w^{*}}}_{\alpha})}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{{(G^{w^{*}})}^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{{w^{*}}}_{\alpha})}\mathbb{Z}^{w^{*}}({\mathcal{T}^{w}_{\alpha}})+\mathbb{Z}^{w^{*}}_{1}(\mathcal{T}^{w}_{\alpha}),$
@@ Appendix D Weighted novelty detection for a general weight function
$\zeta_{\alpha}(t)$ $=1-(1-\pi_{0})\frac{({G_{0}^{w}})^{\prime}G^{w}-{G_{0}^{w}}({G^{w}})^{\prime}}{\mathopen{}\mathclose{{\left(\alpha^{-1}-({G_{{\tiny\mbox{mixt}}}^{w}})^{\prime}(t)}}\right)G_{0}^{w}}$
where $\mathbf{p}(t)$ corresponds to the positive false discovery rate at $t$ (Storey, 2002). Note that in the oracle case $\zeta_{\alpha}(\mathcal{T}^{w^{*}}_{\alpha})=0$, but $\zeta_{\alpha}(\mathcal{T}^{w}_{\alpha})$ is not necessarily zero for a general $w$.
@@
$\sqrt{\tau_{n,m}}\begin{pmatrix}{\widehat{{G}}}_{m,{0}}^{w,(n)}-G_{0}^{w}\\ {\widehat{{G}}}_{m,{1}}^{w,(n)}-G^{w}\end{pmatrix}\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\begin{pmatrix}\frac{\sigma}{\sqrt{\pi_{0}}}\mathbb{U}_{G_{0}^{w}}+({G_{0}^{w}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right)\\ \frac{\sigma}{\sqrt{1-\pi_{0}}}\mathbb{V}_{G^{w}}+({G^{w}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right)\end{pmatrix}\eqqcolon\begin{pmatrix}\mathbb{Z}^{w}_{0}\\ \mathbb{Z}^{w}_{1}\end{pmatrix}$
$\text{ on $\mathopen{}\mathclose{{\left[D(0,1)}}\right]^{2}$}.$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left({\widehat{G}}_{m}^{w,(n)}-{G_{{\tiny\mbox{mixt}}}^{w}}}}\right)\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}$ $\sqrt{\pi_{0}\sigma^{2}}\mathbb{U}_{G_{0}^{w}}+\sqrt{\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\sigma^{2}}\mathbb{V}_{G^{w}}$
$+({{G_{{\tiny\mbox{mixt}}}^{w}}})^{\prime}{\rho^{w}}\sqrt{1-\sigma^{2}}\mathopen{}\mathclose{{\left(\mathbb{W}_{I^{w}}+\mathopen{}\mathclose{{\left[I-I^{w}}}\right]N}}\right)$
$=\pi_{0}\mathbb{Z}^{w}_{0}+\mathopen{}\mathclose{{\left(1-\pi_{0}}}\right)\mathbb{Z}^{w}_{1}\eqqcolon\mathbb{Z}^{w}\text{ on $D(0,1)$}.$
@@
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathcal{T}^{\mathrm{BH}_{\alpha}}({\widehat{G}}_{m}^{w,(n)})-\mathcal{T}^{w}_{\alpha}}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{1}{\frac{1}{\alpha}-({G_{{\tiny\mbox{mixt}}}^{w^{*}}})^{\prime}(\mathcal{T}^{w}_{\alpha})}\mathbb{Z}^{w}({\mathcal{T}^{w}_{\alpha}});$
$\sqrt{\tau_{n,m}}\mathopen{}\mathclose{{\left(\mathrm{FDP}_{m}^{(n)}\mathopen{}\mathclose{{\left({\mathcal{R}}^{w}_{\alpha}}}\right)-\mathbf{p}\mathopen{}\mathclose{{\left(\mathcal{T}^{w}_{\alpha}}}\right)}}\right)$ $\overset{\mathcal{L}}{\underset{\tau_{n,m}\rightarrow+\infty}{\longrightarrow}}\frac{\mathbb{Z}^{w}_{0}({\mathcal{T}^{w}_{\alpha}})}{G_{0}^{w}\mathopen{}\mathclose{{\left(\mathcal{T}^{w}_{\alpha}}}\right)}\mathbf{p}\mathopen{}\mathclose{{\left(\mathcal{T}^{w}_{\alpha}}}\right)\mathopen{}\mathclose{{\left[1-\mathbf{p}\mathopen{}\mathclose{{\left(\mathcal{T}^{w}_{\alpha}}}\right)\zeta_{\alpha}\mathopen{}\mathclose{{\left(\mathcal{T}^{w}_{\alpha}}}\right)}}\right]$